Q1) Suppose the market demand for playing cards is given by the equation

Q = 600 – 100P

Where Q is the no. of decks of cards demand each year and P is the price in Rupee. For a price increase from Rs. 2 to Rs. 3 per deck, what is the price elasticity?















1
Expert's answer
2021-04-24T18:48:37-0400

find Q :

P=2

"Q = 600 - 100\times P=600 - 100\times 2=400"

P=3

"Q = 600 - 100\times P=600 - 100\times 3=300"

price elasticity

"E(DP)=\frac{\%\Delta Q}{\%\Delta P}=\frac{-25}{50}=-0.5"


"\%\Delta Q=\frac{300-400}{400}\times100=-25"


"\%\Delta P=\frac{3-2}{2}\times100=50"